OpenStudy (khally92):

A new cottage is built across the river and 300m downstream from the nearest telephone relay station. the river is 120m wide. In order to wire the cottage for phone service, wire will be laid across the river under water, and along the edge of the river above ground. the cost to lay wire under water is $15 per m and the cost to lay wire above the ground is $10 per m. how much wire should be laid under water to minimize cost? thank you.

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (khally92):

honestly i do not understand... y dont we break it down now?

OpenStudy (anonymous):

From the diagram, x is the length of the wire under water. 300-a is the length of wire above ground. "a" has to be calculated from the right triangle elements. Does this make sense so far?

OpenStudy (khally92):

Yeah it does but how comes x is slant?

OpenStudy (anonymous):

x is slanted to represent a general x position. Remember this is a minimization problem. The possible positions for x is from straight down to directly connected to the cottage across the water.

OpenStudy (khally92):

ok.... to find a now, with two unknows we have square root of x^2-120^2.

OpenStudy (anonymous):

You almost have it. You have to take the square root of x^2-120^2 because a^2 = x^2-120^2. you want "a", not a^2.

OpenStudy (khally92):

x-120=a.

OpenStudy (khally92):

hello are you there?

OpenStudy (anonymous):

\[a=\sqrt{x{}^{\wedge}2-120{}^{\wedge}2} \]I may have to break off. This site/my computor is mis behaving.

OpenStudy (khally92):

when will you come online again please. yeah like, buh if you futher simplify a also = x-120, i.e the square cancels the root. lets go futher please.

OpenStudy (anonymous):

\[a=\sqrt{x^2-120^2} \]cannot be simplified. The cost for the wire above ground is the following:\[10\sqrt{x^2-120^2} \]

OpenStudy (khally92):

yeah am getting it

OpenStudy (anonymous):

15x is the cost of the wire under water.

OpenStudy (khally92):

absolutely. did you multiply the cost by the sides?

OpenStudy (anonymous):

The sum of the above ground and the below water is the total cost to install the wire. \[15x+10\sqrt{x^2-120^2} \]

OpenStudy (khally92):

yeah.

OpenStudy (anonymous):

The derivative of the total cost is:\[15+\frac{10 x}{\sqrt{-14400+x^2}} \]

OpenStudy (anonymous):

Combine the fractions above:\[\frac{5 \left(2 x+3 \sqrt{-14400+x^2}\right)}{\sqrt{-14400+x^2}} \]

OpenStudy (khally92):

combine wat fraction?

OpenStudy (anonymous):

Combine the terms may be a more accurate statement. the result is the following:\[\frac{5 \left(2 x+3 \sqrt{-14400+x^2}\right)}{\sqrt{-14400+x^2}} \]

OpenStudy (anonymous):

Solve \[2 x+3 \sqrt{-14400+x^2}=0 \]for x.

OpenStudy (anonymous):

\[\left(3 \sqrt{-14400+x^2}\right)^2=(-2 x)^2 \]\[9 \left(-14400+x^2\right)=4 x^2 \]\[x=\pm 72 \sqrt{5} \]

OpenStudy (khally92):

you got me confused now.......where did you get 2x +3

OpenStudy (khally92):

what is the derivative of the total cost.

OpenStudy (anonymous):

\[15+\frac{10 x}{\sqrt{-14400+x^2}} \]

OpenStudy (khally92):

the derivative of 10 is 1.

OpenStudy (anonymous):

Total cost is\[15x+10\sqrt{x^2-120^2} \]The derivative of the above is\[15+\frac{10 x}{\sqrt{-14400+x^2}} \]Solve the following for x\[15+\frac{10 x}{\sqrt{-14400+x^2}}=0 \]

OpenStudy (anonymous):

Sorry, but I have to break off now.

OpenStudy (khally92):

am still trying to get it right. thanks bro.

OpenStudy (khally92):

ill get back to you. thanks alot. you are a life saver.